Feather Friends, Inc., distributes a high-quality wooden birdhouse that sells for $120 per unit. Variable expenses are $60.00 per unit, and fixed expenses total $180,000 per year. Its operating results for last year were as follows:

Sales $ 3,120,000 Variable expenses 1,560,000 Contribution margin 1,560,000 Fixed expenses 180,000 Net operating income $ 1,380,000 Required: Answer each question independently based on the original data:

1. What is the product's CM ratio?

2. Use the CM ratio to determine the break-even point in dollar sales.

3. If this year's sales increase by $42,000 and fixed expenses do not change, how much will net operating income increase?

4-a. What is the degree of operating leverage based on last year's sales?

4-b. Assume the president expects this year's sales to increase by 12%. Using the degree of operating leverage from last year, what percentage increase in net operating income will the company realize this year?

5. The sales manager is convinced that a 12% reduction in the selling price, combined with a $64,000 increase in advertising, would increase this year's unit sales by 25%.

5-a. If the sales manager is right, what would be this year's net operating income if his ideas are implemented?

5-b. If the sales manager's ideas are implemented, how much will net operating income increase or decrease over last year?

6. The president does not want to change the selling price. Instead, he wants to increase the sales commission by $2.00 per unit. He thinks that this move, combined with some increase in advertising, would increase this year's sales by 25%. How much could the president increase this year's advertising expense and still earn the same $1,380,000 net operating income as last year?
Answer #1

Contribution Margin Ratio = Contribution Margin/Sales

= 1,560,000/3,120,000

= 50%

Break even point in sales dollars = Fixed costs/Contribution Margin Ratio

= 180,000/50%

= $360,000

Increase in operating income = increase in contribution margin since fixed costs remain same

= 42,000*50%

= $21,000

4-a Degree of Operating leverage = Contribution Margin/Net Operating income

= 1,560,000/1,380,000

= 1.1304

4-b % increase in Income = % increase in sales*Degree of operating leverage

= 12%*1.1304 = 16.956%

i.e. 16.96%

5a. Net Operating income = (105.6 – 60)*26000*1.25 – 180,000-64,000

= $1,238,000

b.Decrease in Net income = -$142,000

6.Desired Operating Income = $1,380,000

Add: Fixed costs other than advertising= 180,000

Total = $1,560,000

Contribution Margin = (120-60-2)*26,000*1.25 = $1,885,000

Increase in advertising expense = 1,885,000-1,560,000 = $325,000